# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-A Auxiliary lemmas
According to the dynamics of $\boldsymbol{x}_{i}^{t}$ in line 6 of Algorithm 1 and the projection inequality, for any $i\in[m]$, $j\neq i$ and $\bar{x}_{(i)}^{t}\in\Omega_{i}$, we have
$\|x_{i}^{t+1}-\bar{x}_{(i)}^{t}\|$ $\textstyle\leq\|x_{i}^{t}+\sum_{j\in\mathcal{N}_{i}}w_{ij}(x_{j(i)}^{t}+\vartheta_{j(i)}^{t}-x_{i}^{t})$
$\textstyle\quad-\lambda_{x}^{t}u_{i}^{t}-\bar{x}_{(i)}^{t}\|,$
$\|x_{j(i)}^{t+1}-\bar{x}_{(i)}^{t}\|$ $\textstyle\leq\|x_{j(i)}^{t}+\sum_{i\in\mathcal{N}_{j}}w_{ji}(x_{i}^{t}+\vartheta_{i}^{t}-x_{j(j)}^{t})$
$\textstyle\quad-\bar{x}_{(i)}^{t}\|,~{}\forall j\neq i.$
The preceding inequalities can be written into a more compact form as follows:
$\textstyle\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|\leq\|((I_{m}+W)\otimes I_{n_{i}})\boldsymbol{x}_{(i)}^{t}$ (24)
$\textstyle\quad+(W^{0}\otimes I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|,$
where $\boldsymbol{u}_{(i)}^{t}$ is given by $\boldsymbol{u}_{(i)}^{t}=\mathrm{col}(\boldsymbol{0}_{n_{i}},\cdots,u_{i},\cdots,\boldsymbol{0}_{n_{i}})$, with the $i$th block being $u_{i}$ and all others being 00, and $W^{0}$ is the same as $W$, except that all of its diagonal entries are zero.
By defining $\tilde{W}=I_{m}+W-\frac{\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{T}}{m}$ and using the relation $\bar{x}_{(i)}^{t}=\frac{\boldsymbol{1}_{m}^{T}\otimes I_{n_{i}}}{m}\boldsymbol{x}_{(i)}^{t}$, we can rewrite (24) as
$\textstyle\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}\leq\|(W^{0}\otimes I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\|^{2}$
$\textstyle\quad+\|(\tilde{W}\otimes I_{n_{i}})(\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t})-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t}\|^{2}$
$\textstyle\quad+2\langle(\tilde{W}\otimes I_{n_{i}})\boldsymbol{x}_{(i)}^{t}-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t},(W^{0}\otimes I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\rangle,$ (25)
where we have subtracted $(\tilde{W}\otimes I_{n_{i}})(\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t})$ from the right hand side of (24) due to $(\tilde{W}\otimes I_{n_{i}})(\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t})=\boldsymbol{0}_{mn_{i}}$.
@@
$\textstyle\|(\tilde{W}\otimes I_{n_{i}})(\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t})-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t}\|^{2}$ (26)
$\textstyle<(1-|\rho_{2}|)\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}+\frac{(\lambda_{x}^{t})^{2}}{|\rho_{2}|}\|\boldsymbol{u}_{(i)}^{t}\|^{2},$
where in the derivation we have used the fact that the inequality $(a+b)^{2}\leq(1+\varepsilon)a^{2}+(1+\varepsilon^{-1})b^{2}$ holds for any scalars $a$, $b$, and $\varepsilon>0$ (by setting $\varepsilon=\frac{|\rho_{2}|}{1-|\rho_{2}|}$, i.e., $1+\varepsilon=\frac{1}{1-|\rho_{2}|}$ and $1+\varepsilon^{-1}=\frac{1}{|\rho_{2}|}$).
@@
$\textstyle\|\boldsymbol{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}\!<\!(1\!-\!|\rho_{2}|)\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}\!+\!\|\boldsymbol{\vartheta}_{(i)}^{t}\|^{2}$
$\textstyle\quad\!+\!\frac{(\lambda_{x}^{t})^{2}}{|\rho_{2}|}\|\boldsymbol{u}_{(i)}^{t}\|^{2}\!+\!2\langle(\tilde{W}\!\otimes\!I_{n_{i}})\boldsymbol{x}_{(i)}^{t}\!-\!\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t},(W^{0}\!\otimes\!I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\rangle.$
Note that when the variables $x_{l(i)}^{t+1}$ are constrained in a convex set $\Omega_{i}$, their average $\bar{x}_{(i)}^{t+1}$ also lies in the set $\Omega_{i}$, implying that the average $\bar{x}_{(i)}^{t+1}$ is the solution to the constrained problem of minimizing $\sum_{l=1}^{m}\|x_{l(i)}^{t+1}-x\|$ over $x\in\Omega_{i}$. Using this and the relation $\sum_{l=1}^{m}\|x_{l(i)}^{t+1}-x\|^{2}=\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes x\|^{2}$, we have $\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}\leq\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}$, which implies
$\textstyle\|\boldsymbol{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\!\otimes\!\bar{x}_{(i)}^{t+1}\|^{2}\!\leq\!(1\!-\!|\rho_{2}|)\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\!\otimes\!\bar{x}_{(i)}^{t}\|^{2}\!+\!\|\boldsymbol{\vartheta}_{(i)}^{t}\|^{2}$
$\textstyle+\!\frac{(\lambda_{x}^{t})^{2}}{|\rho_{2}|}\|\boldsymbol{u}_{(i)}^{t}\|^{2}\!+\!2\langle(\tilde{W}\!\otimes\!I_{n_{i}})\boldsymbol{x}_{(i)}^{t}\!-\!\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t},(W^{0}\!\otimes\!I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\rangle.$
@@
$\textstyle\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}]\!\leq\!(1-|\rho_{2}|)\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$
@@
$\vspace{-0.2em}\textstyle\mathbb{E}[\|\boldsymbol{u}_{(i)}^{t}\|^{2}]\leq 2(\sigma_{f}^{2}+L_{f}^{2})+2(\sigma_{g,1}^{2}+L_{g}^{2})\mathbb{E}[\|\tilde{z}_{i}^{t}\|^{2}],$ (28)
@@
$\textstyle\mathbb{E}[\|\tilde{z}_{i}^{t}\|^{2}]\leq\frac{2}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\hat{z}_{i}^{t+1}-\hat{z}_{i}^{t}\|^{2}]$
$\textstyle+\frac{2}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\bar{\zeta}_{w}^{t}+\lambda_{z}^{t}\nabla_{y}\bar{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\|^{2}]$
$\textstyle\leq\frac{2}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\hat{z}_{i}^{t+1}-\hat{z}_{i}^{t}\|^{2}]+\frac{2r(\sigma_{z}^{t})^{2}}{(\lambda_{z}^{t})^{2}}+2(\sigma_{f}^{2}+L_{f}^{2}),$ (29)
@@
Substituting (29) into (28), we obtain
$\vspace{-0.2em}\textstyle\mathbb{E}[\|\boldsymbol{u}_{(i)}^{t}\|^{2}]\leq\frac{c_{u1}}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\hat{z}_{i}^{t+1}-\hat{z}_{i}^{t}\|^{2}]+\frac{c_{u1}r(\sigma_{z}^{t})^{2}}{(\lambda_{z}^{t})^{2}}+c_{u2},$ (30)
@@
Further incorporating (30) into (27), we arrive at (23). ∎
@@
$\textstyle\quad+c_{y1}(\lambda_{y}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$
@@
where $c_{y1}\!=\!4L_{g}^{2}|\rho_{2}|^{-1}$,$c_{y2}\!=\!c_{y1}c_{u1}$, and $c_{y3}\!=\!2mr+2(\sigma_{g,0}^{2}+d_{g}^{2})mv_{y}^{2}(\lambda_{y}^{0})^{2}|\rho_{2}|^{-1}+4m\sigma_{g,0}^{2}(\lambda_{y}^{1})^{2}|\rho_{2}|^{-1}+\frac{1}{2}c_{y1}n(\lambda_{y}^{1})^{2}\sigma_{x}^{2}+mc_{y1}c_{u1}r(\lambda_{y}^{1})^{2}(\lambda_{x}^{0})^{2}\sigma_{z}^{2}(\lambda_{z}^{0})^{-2}+mc_{y1}c_{u2}(\lambda_{y}^{1})^{2}(\lambda_{x}^{0})^{2}$, with $c_{u1}$ and $c_{u2}$ given in the statement of Lemma 5.
@@
$\boldsymbol{y}^{t+1}=((I_{m}+W)\otimes I_{r})\boldsymbol{y}^{t}+(W^{0}\otimes I_{r})\boldsymbol{\chi}^{t}+\lambda_{y}^{t}\boldsymbol{g}^{t}(\boldsymbol{x}^{t}).\vspace{-0.2em}$ (32)
Using the relations $\bar{y}^{t}=\frac{\boldsymbol{1}_{m}^{T}\otimes I_{r}}{m}\boldsymbol{y}^{t}$ and $\boldsymbol{1}_{m}^{\top}W=\boldsymbol{0}_{m}^{\top}$ yields
$\bar{y}^{t+1}=\bar{y}^{t}+\frac{\boldsymbol{1}_{m}^{T}\otimes I_{r}}{m}(W^{0}\otimes I_{r})\boldsymbol{\chi}^{t}+\lambda_{y}^{t}\frac{\boldsymbol{1}_{m}^{T}\otimes I_{r}}{m}\boldsymbol{g}^{t}(\boldsymbol{x}^{t}).\vspace{-0.2em}$ (33)
Combining (32) and (33), we have
@@
Recalling the definition $\boldsymbol{\hat{y}}^{t}=\boldsymbol{y}^{t}-\boldsymbol{1}_{m}\otimes\bar{y}^{t}$, we have
$\textstyle\mathbb{E}[\|\boldsymbol{\hat{y}}^{t+1}-\boldsymbol{\hat{y}}^{t}\|^{2}]\leq\mathbb{E}[\|(\tilde{W}\otimes I_{r})(\boldsymbol{\hat{y}}^{t}-\boldsymbol{\hat{y}}^{t-1})$ (35)
$\textstyle\quad+(\Gamma\otimes I_{r})(\lambda_{y}^{t}\boldsymbol{g}^{t}(\boldsymbol{x}^{t})-\lambda_{y}^{t-1}\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1}))\|^{2}]$
$\textstyle\quad+\mathbb{E}[\|\boldsymbol{\chi}^{t}-\boldsymbol{\chi}^{t-1}\|^{2}].$
Using an argument similar to the derivation of (27) yields
@@
By defining $\vartheta_{w(i)}^{t}\triangleq\sum_{j\in\mathcal{N}_{i}}w_{ij}\vartheta_{j(i)}^{t}$ and using the projection inequality, the last term on the right hand side of (38) satisfies
$\textstyle\mathbb{E}[\|x_{i}^{t}-x_{i}^{t-1}\|^{2}]\leq 2\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$ (39)
$\textstyle\quad+2(\lambda_{x}^{t-1})^{2}\mathbb{E}[\|u_{i}^{t-1}\|^{2}]\!+\!n_{i}(\sigma_{x}^{t-1})^{2}.$
According to the relation $\mathbb{E}[\|u_{i}^{t}\|^{2}]=\mathbb{E}[\|\boldsymbol{u}_{(i)}^{t}\|^{2}]$, we incorporate (30) into (39) to obtain
$\textstyle\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{t-1}\|^{2}]\leq 2\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$
@@
$\textstyle 2(\lambda_{y}^{t})^{2}\mathbb{E}[\|\boldsymbol{g}^{t}(\boldsymbol{x}^{t})-\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})\|^{2}]\leq 4m\sigma_{g,0}^{2}(\lambda_{y}^{t})^{2}t^{-1}$ (41)
$\textstyle\quad+4L_{g}^{2}(\lambda_{y}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$
$\textstyle\quad+\!\frac{4L_{g}^{2}c_{u1}(\lambda_{y}^{t})^{2}(\lambda_{x}^{t-1})^{2}}{(\lambda_{z}^{t-1})^{2}}\mathbb{E}[\|\boldsymbol{\hat{z}}^{t}\!-\!\boldsymbol{\hat{z}}^{t-1}\|^{2}]\!+\!2nL_{g}^{2}(\lambda_{y}^{t})^{2}(\sigma_{x}^{t-1})^{2}$
$\textstyle\quad+\!\frac{4mL_{g}^{2}c_{u1}r(\lambda_{y}^{t})^{2}(\lambda_{x}^{t-1})^{2}(\sigma_{z}^{t-1})^{2}}{(\lambda_{z}^{t-1})^{2}}\!+\!4mL_{g}^{2}c_{u2}(\lambda_{y}^{t})^{2}(\lambda_{x}^{t-1})^{2}.$
Next, we characterize the second term on the right-hand side of (37). Using the mean-value theorem, we have
$\textstyle(\lambda_{y}^{t}-\lambda_{y}^{t-1})^{2}\!=\!\frac{\lambda_{y}^{0}}{(t+1)^{v_{y}}}-\frac{\lambda_{y}^{0}}{t^{v_{y}}}\!=\!\left(\frac{v_{y}\lambda_{y}^{0}}{\varepsilon^{1+v_{y}}}\right)^{2}\!\leq\!\left(\frac{v_{y}\lambda_{y}^{0}}{t^{1+v_{y}}}\right)^{2},$ (42)
for some $\varepsilon\in(t,t+1)$. Therefore, the second term on the right hand side of (37) satisfies
@@
where we have used the fact that the optimization variable $x_{i}^{t}$ is always constrained in the compact set $\Omega_{i}$, i.e., $\max_{i\in[m]}\sup_{x_{i}\in\Omega_{i}}\{\|g_{i}(x_{i})\|\}\leq d_{g}$ holds for some $d_{g}>0$.
Substituting (41) and (43) into (37), and further substituting (37) into (36), we arrive at (31). ∎
@@
$\textstyle+c_{z1}(\lambda_{z}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$
@@
The definitions of $\boldsymbol{\tilde{y}}^{t}$ and $\boldsymbol{\hat{y}}^{t}$ imply
$\textstyle\|\boldsymbol{\tilde{y}}^{t}-\boldsymbol{\tilde{y}}^{t-1}\|^{2}\leq\frac{4}{(\lambda_{y}^{t})^{2}}\left(\|\boldsymbol{\hat{y}}^{t+1}-\boldsymbol{\hat{y}}^{t}\|^{2}+m\|\bar{y}^{t+1}-\bar{y}^{t}\|^{2}\right)$
$\textstyle\quad+\!\frac{4}{(\lambda_{y}^{t-1})^{2}}\left(\|\boldsymbol{\hat{y}}^{t}-\boldsymbol{\hat{y}}^{t-1}\|^{2}+m\|\bar{y}^{t}-\bar{y}^{t-1}\|^{2}\right).$ (47)
Incorporating (33) into (47) and using an argument similar to the derivation of (43), we obtain

### v3

@@ VII-A Auxiliary lemmas
According to the dynamics of $\mathbf{x}_{i}^{t}$ and the projection inequality, for any $i\in[m]$, $j\neq i$ and $\bar{\mathrm{x}}_{(i)}^{t}\in\Omega_{i}$, we have
$\|x_{i}^{t+1}-\bar{\mathrm{x}}_{(i)}^{t}\|$ $\textstyle\leq\|x_{i}^{t}+\sum_{j\in\mathcal{N}_{i}}w_{ij}(\mathrm{x}_{j(i)}^{t}+\vartheta_{j(i)}^{t}-x_{i}^{t})$
$\textstyle\quad-\lambda_{x}^{t}u_{i}^{t}-\bar{\mathrm{x}}_{(i)}^{t}\|,$
$\|\mathrm{x}_{j(i)}^{t+1}-\bar{\mathrm{x}}_{(i)}^{t}\|$ $\textstyle\leq\|\mathrm{x}_{j(i)}^{t}+\sum_{l\in\mathcal{N}_{j}}w_{jl}(\mathrm{x}_{l(i)}^{t}+\vartheta_{l(i)}^{t}-\mathrm{x}_{j(i)}^{t})$
$\textstyle\quad-\bar{\mathrm{x}}_{(i)}^{t}\|,~\forall j\neq i.$
The compact form of the preceding inequalities is
$\textstyle\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|\leq\|((I_{m}+W)\otimes I_{n_{i}})\mathbf{x}_{(i)}^{t}$ (25)
$\textstyle\quad+(W^{0}\otimes I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|,$
where $\boldsymbol{u}_{(i)}^{t}$ is given by $\boldsymbol{u}_{(i)}^{t}=\mathrm{col}(\boldsymbol{0}_{n_{i}},\cdots,u_{i},\cdots,\boldsymbol{0}_{n_{i}})$, with the $i$th block being $u_{i}$ and all others being 0, and $W^{0}$ is the same as $W$, except that all of its diagonal entries are zero.
By defining $\tilde{W}=I_{m}+W-\frac{\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{T}}{m}$ and using the relation $\bar{\mathrm{x}}_{(i)}^{t}=\frac{\boldsymbol{1}_{m}^{T}\otimes I_{n_{i}}}{m}\mathbf{x}_{(i)}^{t}$, we can rewrite (25) as
$\textstyle\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}\leq\|(W^{0}\otimes I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\|^{2}$
$\textstyle\quad+\|(\tilde{W}\otimes I_{n_{i}})(\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t})-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t}\|^{2}$
$\textstyle\quad+2\langle(\tilde{W}\otimes I_{n_{i}})\mathbf{x}_{(i)}^{t}-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t},(W^{0}\otimes I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\rangle,$ (26)
where we have subtracted $(\tilde{W}\otimes I_{n_{i}})(\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t})$ from the right hand side of (25) due to $(\tilde{W}\otimes I_{n_{i}})(\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t})=\boldsymbol{0}_{mn_{i}}$.
@@
$\textstyle\|(\tilde{W}\otimes I_{n_{i}})(\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t})-\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t}\|^{2}$ (27)
$\textstyle<(1-|\rho_{2}|)\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}+\frac{(\lambda_{x}^{t})^{2}}{|\rho_{2}|}\|\boldsymbol{u}_{(i)}^{t}\|^{2},$
where in the derivation we have used the inequality $(a+b)^{2}\leq(1+\varepsilon)a^{2}+(1+\varepsilon^{-1})b^{2}$ for any scalars $a$, $b$, and $\varepsilon>0$ (by setting $\varepsilon=\frac{|\rho_{2}|}{1-|\rho_{2}|}$, i.e., $1+\varepsilon=\frac{1}{1-|\rho_{2}|}$ and $1+\varepsilon^{-1}=\frac{1}{|\rho_{2}|}$).
@@
$\textstyle\|\mathbf{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}\!<\!(1\!-\!|\rho_{2}|)\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}\!+\!\|\boldsymbol{\vartheta}_{(i)}^{t}\|^{2}$
$\textstyle\quad\!+\!\frac{(\lambda_{x}^{t})^{2}}{|\rho_{2}|}\|\boldsymbol{u}_{(i)}^{t}\|^{2}\!+\!2\langle(\tilde{W}\!\otimes\!I_{n_{i}})\mathbf{x}_{(i)}^{t}\!-\!\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t},(W^{0}\!\otimes\!I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\rangle.$
Note that when the variables $\mathrm{x}_{l(i)}^{t+1}$ are constrained in a convex set $\Omega_{i}$, their average $\bar{\mathrm{x}}_{(i)}^{t+1}$ also lies in the set $\Omega_{i}$, implying that the average $\bar{\mathrm{x}}_{(i)}^{t+1}$ is the solution to the constrained problem of minimizing $\sum_{l=1}^{m}\|\mathrm{x}_{l(i)}^{t+1}-x\|$ over $x\in\Omega_{i}$. Using this and the relation $\sum_{l=1}^{m}\|\mathrm{x}_{l(i)}^{t+1}-x\|^{2}=\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes x\|^{2}$, we have $\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}\leq\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}$, which implies
$\textstyle\|\mathbf{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\!\otimes\!\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}\!\leq\!(1\!-\!|\rho_{2}|)\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\!\otimes\!\bar{\mathrm{x}}_{(i)}^{t}\|^{2}\!+\!\|\boldsymbol{\vartheta}_{(i)}^{t}\|^{2}$
$\textstyle+\!\frac{(\lambda_{x}^{t})^{2}}{|\rho_{2}|}\|\boldsymbol{u}_{(i)}^{t}\|^{2}\!+\!2\langle(\tilde{W}\!\otimes\!I_{n_{i}})\mathbf{x}_{(i)}^{t}\!-\!\lambda_{x}^{t}\boldsymbol{u}_{(i)}^{t},(W^{0}\!\otimes\!I_{n_{i}})\boldsymbol{\vartheta}_{(i)}^{t}\rangle.$
@@
$\textstyle\mathbb{E}[\|\mathbf{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}]\!\leq\!(1-|\rho_{2}|)\mathbb{E}[\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]$
@@
$\vskip-1.99997pt\textstyle\mathbb{E}[\|\boldsymbol{u}_{(i)}^{t}\|^{2}]\leq 2(\sigma_{f}^{2}+L_{f}^{2})+2(\sigma_{g,1}^{2}+L_{g}^{2})\mathbb{E}[\|\tilde{z}_{i}^{t}\|^{2}],$ (29)
@@
$\mathbb{E}[\|\tilde{z}_{i}^{t}\|^{2}]\leq\frac{2}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\hat{z}_{i}^{t+1}-\hat{z}_{i}^{t}\|^{2}]+\frac{2r(\sigma_{z}^{t})^{2}}{(\lambda_{z}^{t})^{2}}+2(\sigma_{f}^{2}+L_{f}^{2}),$
@@
Substituting the preceding inequality into (29), we obtain
$\mathbb{E}[\|\boldsymbol{u}_{(i)}^{t}\|^{2}]\leq\frac{c_{u1}}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\hat{z}_{i}^{t+1}-\hat{z}_{i}^{t}\|^{2}]+\frac{c_{u1}r(\sigma_{z}^{t})^{2}}{(\lambda_{z}^{t})^{2}}+c_{u2},$ (30)
@@
Further incorporating (30) into (28), we arrive at (24).
@@
$\textstyle\quad+c_{y1}(\lambda_{y}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t-1}\|^{2}]$
@@
where $c_{y1}\!=\!4L_{g}^{2}|\rho_{2}|^{-1}$$c_{y2}\!=\!c_{y1}c_{u1}$, and $c_{y3}\!=\!2mr+2(\sigma_{g,0}^{2}+d_{g}^{2})mv_{y}^{2}(\lambda_{y}^{0})^{2}|\rho_{2}|^{-1}+4m\sigma_{g,0}^{2}(\lambda_{y}^{1})^{2}|\rho_{2}|^{-1}+\frac{1}{2}c_{y1}n(\lambda_{y}^{1})^{2}\sigma_{x}^{2}+mc_{y1}c_{u1}r(\lambda_{y}^{1})^{2}(\lambda_{x}^{0})^{2}\sigma_{z}^{2}(\lambda_{z}^{0})^{-2}+mc_{y1}c_{u2}(\lambda_{y}^{1})^{2}(\lambda_{x}^{0})^{2}$, with $c_{u1}$ and $c_{u2}$ given in the statement of Lemma 5.
@@
Recalling the definition $\boldsymbol{\hat{y}}^{t}=\boldsymbol{y}^{t}-\boldsymbol{1}_{m}\otimes\bar{y}^{t}$ and using an argument similar to the derivation of (28) yield
@@
By defining $\vartheta_{w(i)}^{t}\triangleq\sum_{j\in\mathcal{N}_{i}}w_{ij}\vartheta_{j(i)}^{t}$ and and using (30), the last term on the right hand side of (35) satisfies
$\textstyle\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{t-1}\|^{2}]\leq 2\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t-1}\|^{2}]$
@@
$\textstyle 2(\lambda_{y}^{t})^{2}\mathbb{E}[\|\boldsymbol{g}^{t}(\boldsymbol{x}^{t})-\boldsymbol{g}^{t-1}(\boldsymbol{x}^{t-1})\|^{2}]\leq 4m\sigma_{g,0}^{2}(\lambda_{y}^{t})^{2}t^{-1}$
$\textstyle\!\!+4L_{g}^{2}(\lambda_{y}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t-1}\|^{2}]$
$\textstyle\!\!+\frac{4L_{g}^{2}c_{u1}(\lambda_{y}^{t})^{2}(\lambda_{x}^{t-1})^{2}}{(\lambda_{z}^{t-1})^{2}}\mathbb{E}[\|\boldsymbol{\hat{z}}^{t}\!-\!\boldsymbol{\hat{z}}^{t-1}\|^{2}]\!+\!2nL_{g}^{2}(\lambda_{y}^{t})^{2}(\sigma_{x}^{t-1})^{2}$
$\textstyle\!\!+\left(\frac{4mL_{g}^{2}c_{u1}r(\sigma_{z}^{t-1})^{2}}{(\lambda_{z}^{t-1})^{2}}\!+\!4mL_{g}^{2}c_{u2}\right)(\lambda_{y}^{t})^{2}(\lambda_{x}^{t-1})^{2}.$ (37)
Using the relation $(\lambda_{y}^{t}-\lambda_{y}^{t-1})^{2}\leq\left(\frac{v_{y}\lambda_{y}^{0}}{t^{1+v_{y}}}\right)^{2}$ for some $\varepsilon\in(t,t+1)$, the second term on the right hand side of (34) satisfies
@@
where we have used $\max_{i\in[m]}\sup_{x_{i}\in\Omega_{i}}\{\|g_{i}(x_{i})\|\}\leq d_{g}$ for some $d_{g}>0$. Substituting (37) and (38) into (34), and further substituting (34) into (33), we arrive at (31). ∎
@@
$\textstyle+c_{z1}(\lambda_{z}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t-1}\|^{2}]$
@@
Using an argument similar to the derivation of (38) yields